# Towards Real-time Adaptive Anisotropic Image-to-mesh Conversion for Vascular Flow Simulations

### v1

@@ Abstract
Presented is a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries within the Computational Fluid Dynamics (CFD) modeling process. The proposed method combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies quality, fidelity, and smoothness requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies real-time requirements, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models. The proposed approach is tested with two brain aneurysm cases and is shown to satisfy all the aforementioned requirements. The next steps are to fully parallelize the remaining components of the pipeline to maximize potential performance and to test its integration within a CFD vascular flow simulation. Just as the parallel anisotropic adaptation procedure was tested within aerospace CFD simulations using CAD models, the method is expected to provide accurate results for CFD vascular flow simulations in real-time when executed on multicore cc-NUMA architectures.
1 Introduction.
@@
Interventional radiology involves minimally invasive, catheter-based, deployment of several vascular diseases throughout the body. This field has seen an explosion of device technology over the past decade ranging from stent-grafts for aortic aneurysms to valve replacement for diseased cardiac valves. Neuro-interventional radiology, which deals with diseases of the cerebral vasculature, has expanded from the treatment of brain aneurysms to removal of blood clots in large vessel occlusion strokes to embolization of the middle meningeal artery for treatment of chronic subdural hematomas to venous sinus stenting for treatment of idiopathic intracranial hypertension and pulsatile tinnitus. As hemodynamics are an intrinsic factor of vascular disease, there has been a concomitant increase in the number of studies using Computational Fluid Dynamics (CFD) and/or fluid-structure interactions to better understand disease prognosis or treatment outcomes after deployment of the pertinent medical device. For example, the transport of cardiogenic emboli causing ischemic stroke can be evaluated using imaging scans of the entire vascular branching network in each patient [51]. The recent introduction of miniaturized optical coherence tomography allows for high-resolution imaging of medical device features as well as the vascular wall in any given patient [53], which in turn facilitates high-resolution computational studies of the pathology. CFD may be able to provide clarity in the device treatment of cerebral venous disease, which has irregular morphologies and complex flow patterns [62].
@@
Brain aneurysms (such as that shown in fig. 1) are abnormal focal dilations of intracranial arteries which when left untreated may rupture, resulting in fatal outcomes for patients. Aneurysms have been treated by neuro-interventional devices for nearly three decades, and hence, numerous CFD studies have been conducted on this disease. High resolution CFD simulations provide the opportunity to study hemodynamics in arteries, which can assist clinicians in determining an aneurysm’s potential to rupture and possible methods of treatment (such as the design of neurovascular stents) [52, 13, 17, 12, 36, 46]. Brain aneurysms that form at vascular bifurcations may be subjected to flow instabilities that can only be captured by high-fidelity CFD [41]. Stents used to treat brain aneurysms have wire sizes that can be as small as 25-30 microns with aneurysm sizes ranging from millimeters to centimeters. CFD studies on the effect of such devices (including overlapping devices) on intraneurysmal hemodynamics involve generation of volumetric meshes with elements at multiple sizes [24]. Processing patient-specific medical imaging data of cerebral aneurysms involves the discretization of these complex bodies in the form of smoothed free-form anisotropic body-conforming meshes. Mesh generation methods utilized in commercially available products work well when processing patient-specific cerebral aneurysm geometries; however, these methods do not always satisfy the time constraints of high resolution CFD simulations.
Figure 1: A) an angiogram showing a large $\sim$25 mm aneurysm on the right internal carotid artery; B) vasculature segmented in SimVascular [3], lofted in Rhinoceros [2], and imported into SolidWorks [4]; C) flow diverter stent simulated in Matlab with VMTK [9] input; D) vasculature and stent bodies imported into ADINA [1] for mesh generation (also appears in [25]).
@@
These types of medical simulations can be classified into two categories: interactive and predictive. Interactive simulations require real-time mesh generation which can be used to allow surgical residents to practice performing a procedure within a virtual environment. Real-time mesh generation is also required for image registration applications, as they provide guidance to surgeons during operations [21]. Predictive simulations predict and optimize the outcome of an intervention using patient-specific pre-operative image data. Although predictive simulations don’t necessarily require real-time mesh generation, both types of simulations are time-sensitive given their impact on a patient’s treatment and associated costs.
@@
This paper presents a path towards a fast and robust adaptive anisotropic mesh generation method that is designed to help streamline the discretization of complex vascular geometries and Computer-Aided Design (CAD)-based stents [25, 40] within the CFD modeling process. Several requirements should be addressed with regards to medical image-to-mesh (I2M) conversion: real-time mesh generation, robustness, fidelity, quality, and smoothness. Robustness concerns the software’s ability to process different input types such as CAD models (for example, of stents) and patient-specific medical images. Utilizing both is essential when running a CFD vascular flow simulation meant to aid in determining potential methods of treatment [40, 26, 25]. Fidelity measures the degree to which the mesh surface aligns with an image boundary. Quality (determined by the shape and size of mesh elements) and smoothness affect the accuracy of solutions for CFD simulations [10, 32]. Each of these requirements are independently challenging problems. The proposed approach addresses the robustness, fidelity, quality, and (to a degree) smoothness requirements (providing a smoothness of $C^{0}$, as opposed to other available real-time I2M methods that do not address smoothness [31]). Given that the image-to-mesh conversion and boundary layer generation steps are sequential, the proposed approach only addresses the real-time requirement with its parallel adaptive anisotropic mesh generation. After this feasibility study, the next steps will be to fully parallelize all components of the proposed approach in order to maximize potential performance.
@@
The proposed method involves several steps: (1) utilize a method known as CBC3D [27] to discretize a segmented image using a Body-Centered Cubic lattice of high-quality tetrahedra and deform the generated mesh surfaces to their corresponding tissue boundaries to improve fidelity while maintaining quality, (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid outside the high-fidelity CBC3D mesh surface using a method known as Advancing Front Local Reconnection (AFLR) [48], and (4) utilize a multicore cc-NUMA-based mesh adaptation method known as CDT3D [64] to generate an anisotropic mesh volume using the metric tensor field and CBC3D isotropic mesh as a background mesh. We evaluate this approach using two aneurysm cases (a carotid cavernous aneurysm and middle cerebral artery bifurcation aneurysm) obtained from the Cerebrovascular Center for Research at Stony Brook University under Institutional Review Board approval. We show that the proposed approach meets the aforementioned I2M conversion requirements. Given that CDT3D also exhibited excellent performance and accurately captured features of underlying aerospace simulations when coupled with a solver (processing CAD data such as that shown in fig. 2), the end goal is to utilize this method within a vascular flow simulation.
Figure 2: A CDT3D mesh of an aerospace geometry generated from CAD data and its associated solution obtained from a CFD simulation are shown (also appears in [64]).
2 Related Work.
@@
There are numerous methods that process medical data as input to create meshes. The type of meshes generated must be taken into account when considering the target application. Meshes can be classified into two categories - structured and unstructured. While some structured mesh generation methods have been shown to provide accurate results within CFD simulations [54, 38, 59], these types of methods can sometimes fail to capture features within complex geometries [11]. Unstructured meshes can be further classified into two categories - isotropic and anisotropic. Isotropic mesh generation methods are well suited for high curvature geometries. Anisotropic meshes contain high aspect ratio elements that include directional information (e.g., hemodynamics). Additionally, they may contain fewer elements in some cases compared to isotropic meshes generated from the same image. A mesh should ideally be as small as possible to reduce the memory and CPU requirements for the solver. These benefits make anisotropic meshes suitable for blood flow problems.
@@
There are many methods that convert 3D images into 3D meshes. For a comprehensive overview of these methods, see [68]. Many 3D isotropic methods utilize Delaunay refinement techniques [14, 56, 34, 60, 16, 58, 18, 31, 63, 30]. A challenge with Delaunay refinement, which still remains an open problem, is that almost flat tetrahedra (slivers) can survive known heuristics designed to remove them [14, 18, 20]. These low quality elements introduce error when processed by a solver. Additional I2M methods utilize lattice space-tree (octree) decomposition techniques [15, 70, 69, 42, 19, 67, 71, 44, 37, 45, 57, 50] to provide meshes with adaptively-sized, high quality elements. Some also utilize the Dual Contouring technique to improve element quality [33, 71, 44, 37, 45]. Others warp the surfaces of the lattice to the boundaries of the object using either a mass-spring system, an FE constitutive model, or an optimization scheme [50, 46, 34]. Some methods focus on processing surface geometries of medical data to create structured Cartesian cut-cell meshes [55] and isotropic tetrahedral meshes [39]. [29] evaluates several of these sequential isotropic-based methods’ suitability for surgical simulations and non-rigid registration applications. There are methods which utilize parallel mesh generation techniques (with either CPUs [45, 54, 38, 59] or GPUs [45]). However, [54, 38, 59] all focus on generating Cartesian cut-cell meshes while [45] focuses on generating isotropic tetrahedral meshes, as with all the aforementioned Delaunay and lattice-based methods. [37] utilizes an anisotropic Giaquinta–Hildebrandt operator-based geometric flow method to smooth mesh surfaces while preserving surface and volume features. However, this smoothing is applied to isotropic meshes.
@@
Some methods focus on converting medical images into anisotropic meshes. Although potentially useful for blood flow simulations, [22] focuses on using an anisotropic mesh adaptation procedure to construct a high-fidelity virtual model of a fractured pelvis, and then to build a physical replica (by means of additive manufacturing techniques) to represent the anatomical object from a segmented image. [16] generates high-quality adaptive hybrid meshes for fluid-solid geometries within cardiac environments. Following segmentation, surfaces between heart tissue and blood (and other non-cardiovascular tissue/fluids) are extracted and adapted. Anisotropic tetrahedra are generated for the interior of heart tissue, a prismatic boundary layer is generated for the interface between tissue and blood, and TetGen (an isotropic boundary-constrained Delaunay-based method [61]) is used to tetrahedralize the remainder of the blood region. Boundary layers in meshes are useful when calculating metrics such as pressure or wall shear stress (WSS) caused by blood flow [52, 16]. Although the method generated high-quality hybrid meshes, the Delaunay-based isotropic elements in fluid flow regions (such as in the coronary arteries) were identified as a drawback, with anisotropy suggested as an improvement. Additionally, Delaunay tetrahedralization limits the quality of the tetrahedral core for complex biological geometries [16] and can fail to terminate for certain cases [27]. Another method was developed to model cerebral aneurysms for fluid-structure interaction simulations [66]. The surface mesh of a healthy blood vessel is deformed onto an aneurysm model, simulating the formation of the aneurysm. While the simulation results demonstrated the benefit of utilizing anisotropic meshes rather than isotropic meshes, the mesh generation method is sequential, as with all the aforementioned adaptive anisotropic methods.
The proposed approach combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies the quality, fidelity, and smoothness ($C^{0}$) requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which satisfies the real-time requirement, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models (such as that shown in fig. 2 [64]). To the best of our knowledge, there is no single method that provides all of the above.
3 Background.
3.1 Image-to-mesh Conversion.

### v3

# Near Real-time Adaptive Isotropic and Anisotropic Image-to-mesh Conversion for Numerical Simulations Involving Cerebral Aneurysms

@@ Abstract
Presented are two techniques that are designed to help streamline the discretization of complex vascular geometries within the numerical modeling process. The first method integrates multiple software tools into a single pipeline which can generate adaptive anisotropic meshes from segmented medical images. The pipeline is shown to satisfy quality, fidelity, smoothness, and robustness requirements while providing near real-time performance for medical image-to-mesh conversion. The second method approximates a user-defined sizing function to generate adaptive isotropic meshes of good quality and fidelity in real-time. Tested with two brain aneurysm cases and utilizing up to 96 CPU cores within a single, multicore node on Purdue University’s Anvil supercomputer, the parallel adaptive anisotropic meshing method utilizes a hierarchical load balancing model (designed for large, cc-NUMA shared memory architectures) and contains an optimized local reconnection operation that performs three times faster than its original implementation from previous studies. The adaptive isotropic method is shown to generate a mesh of up to approximately 50 million elements in less than a minute while the adaptive anisotropic method is shown to generate approximately the same number of elements in about 5 minutes.
keywords:
image-to-mesh conversion , high-performance computing , Anvil supercomputer , aneurysms , numerical simulations
journal: Journal of Parallel and Distributed Computing
[…]
1 Introduction
@@
Interventional radiology involves minimally invasive, catheter-based, deployment of medical devices to treat several vascular diseases throughout the body. This field has seen an explosion of device technology over the past decade ranging from stent-grafts for aortic aneurysms to valve replacement for diseased cardiac valves. Neuro-interventional radiology, which deals with diseases of the cerebral vasculature, has expanded from the treatment of brain aneurysms to removal of blood clots in large vessel occlusion strokes to embolization of the middle meningeal artery for treatment of chronic subdural hematomas to venous sinus stenting for treatment of idiopathic intracranial hypertension and pulsatile tinnitus. As hemodynamics are an intrinsic factor of vascular disease, there has been a concomitant increase in the number of studies using Computational Fluid Dynamics (CFD) and/or fluid-structure interactions to better understand disease prognosis or treatment outcomes after deployment of the pertinent medical device. For example, the transport of cardiogenic emboli causing ischemic stroke can be evaluated using imaging scans of the entire vascular branching network in each patient [RoleOfCircle2018]. The recent introduction of miniaturized optical coherence tomography allows for high-resolution imaging of medical device features as well as the vascular wall in any given patient [VolumetricMicroscopy2024], which in turn facilitates high-resolution computational studies of the pathology. CFD may be able to provide clarity in the device treatment of cerebral venous disease, which has irregular morphologies and complex flow patterns [SoVein2023].
@@
Brain aneurysms are abnormal focal dilations of intracranial arteries which when left untreated may rupture, resulting in fatal outcomes for patients. Aneurysms have been treated by neuro-interventional devices for nearly three decades, and hence, numerous numerical modeling studies have been conducted on this disease. High resolution CFD simulations provide the opportunity to study hemodynamics in arteries, which can assist clinicians in determining an aneurysm’s potential to rupture and possible methods of treatment (such as the design of neurovascular stents) [CFDAneurysmRupture2019, HemodynamicReview2019, EfficientSimulation2005, ComputationalVascular2010, OptimizationOfCardiovascular2012, CBC3DYixun]. Brain aneurysms that form at vascular bifurcations may be subjected to flow instabilities that can only be captured by high-fidelity CFD [OnThePrevalence2021]. Stents used to treat brain aneurysms have wire sizes that can be as small as 25-30 microns with aneurysm sizes ranging from millimeters to centimeters. CFD studies on the effect of such devices (including overlapping devices) on intraneurysmal hemodynamics involve generation of volumetric meshes with elements at multiple sizes [CompactingSingle2017]. Processing patient-specific medical imaging data of cerebral aneurysms involves the discretization of these complex bodies in the form of smoothed free-form anisotropic body-conforming meshes. Mesh generation methods utilized in commercially available products work well when processing patient-specific cerebral aneurysm geometries; however, these methods do not always satisfy the time constraints of high resolution CFD simulations.
@@
These types of medical simulations can be classified into two categories: interactive and predictive. Interactive simulations require real-time mesh generation which can be used to allow surgical residents to practice performing a procedure within a virtual environment. Real-time mesh generation is also required for image registration applications, as they provide guidance to surgeons during operations [ComparisonPhysicsBased2023]. Predictive simulations predict and optimize the outcome of an intervention using patient-specific pre-operative image data. Although predictive simulations don’t necessarily require real-time mesh generation, both types of simulations are time-sensitive given their impact on a patient’s treatment and associated costs.
@@
This paper presents two techniques that are designed to help streamline the discretization of complex vascular geometries within the numerical modeling process. Several requirements should be addressed with regards to medical image-to-mesh (I2M) conversion: real-time mesh generation, robustness, fidelity, quality, and smoothness. Robustness concerns the software’s ability to process different input types such as CAD models (for example, of stents) and patient-specific medical images. Utilizing both is essential when running a medical simulation meant to aid in determining potential methods of treatment [MelinaNumericalSimulation2016, InVitro2015, HighFidelity2015]. Fidelity measures the degree to which the mesh surface aligns with an image boundary. Quality (determined by the shape and size of mesh elements) and smoothness affect the accuracy of solutions for CFD simulations [meshEvaluationNRR]. Each of these requirements are independently challenging problems. The first proposed method integrates multiple software tools into a single pipeline to address the robustness, fidelity, quality, and (to a degree) smoothness requirements (providing a smoothness of $C^{0}$, as opposed to other available real-time I2M methods that do not address smoothness [foteinos2014high]). Given that the image-to-mesh conversion and boundary layer generation steps are sequential, the pipeline only addresses the real-time requirement with its parallel adaptive anisotropic mesh generation. After this feasibility study, the next steps will be to fully parallelize all components of the I2M pipeline in order to maximize potential performance. The I2M pipeline involves several steps: (1) utilize a method known as CBC3D [EwCCBC3D] to discretize a segmented image using a Body-Centered Cubic lattice of high-quality tetrahedra and deform the generated mesh surfaces to their corresponding tissue boundaries to improve fidelity while maintaining quality, (2) construct an anisotropic metric tensor field, (3) generate a boundary layer grid from the high-fidelity CBC3D mesh surface using a method known as Advancing Front Local Reconnection (AFLR) [AFLRBoundaryLayer2013], and (4) utilize a multicore cc-NUMA-based mesh adaptation method known as CDT3D [EwCCDT3D] to generate an anisotropic mesh volume using the metric tensor field and CBC3D isotropic mesh as a background mesh. To help satisfy real-time requirements, CDT3D utilizes an optimized local reconnection operation and leverages a hierarchical model to balance workloads between threads.
@@
The second method extends the functionality of a Delaunay-based image-to-mesh conversion software called PODM [foteinos2014high]. The method can now utilize a user-defined sizing function to adapt volume elements when generating an isotropic mesh from a segmented image, while satisfying fidelity, quality, and real-time requirements. We evaluate these methods using two aneurysm cases (a carotid cavernous aneurysm and middle cerebral artery bifurcation aneurysm) obtained from the Cerebrovascular Center for Research at Stony Brook University under Institutional Review Board approval. We show that the proposed approaches meet the aforementioned I2M conversion requirements. Additionally, when utilizing 96 CPU cores on a single, multicore node on Purdue University’s Anvil supercomputer [Anvil2022], the adaptive isotropic method generates about 50 million elements in less than a minute while the adaptive anisotropic CDT3D method generates approximately the same amount in about 5 minutes. Given that CDT3D also exhibited excellent performance and accurately captured features of underlying aerospace simulations when coupled with a solver (also processing CAD data to perform adaptation [EwCCDT3D]), the end goal is to utilize this method within a vascular flow simulation involving stents. This is however outside the scope of this paper and we first test the feasibility of utilizing the methods to help satisfy the aforementioned image-to-mesh conversion requirements.
Ultimately, this paper contributes the following:
  1. 1.
    an image-to-mesh conversion pipeline for adaptive anisotropic mesh generation that satisfies robustness, quality, fidelity, smoothness, and near real-time medical I2M conversion requirements,
  2. 2.
    a hierarchical load balancing model (designed to target large cc-NUMA shared memory architectures) for speculative execution applied to an adaptive anisotropic mesh generation method that only utilized 40 CPU cores for aerospace cases in earlier studies [EwCCDT3D, TsolakisEvaluation] and is now tested up to 96 for medical cases,
  3. 3.
    an optimized local reconnection algorithm for anisotropic adaptation, that when compared to its implementation in earlier studies [EwCCDT3D, TsolakisEvaluation], is now three times faster,
  4. 4.
    and a new sizing function approximation technique to generate adaptive isotropic meshes from segmented medical images.
Section 2 presents an overview of related work. Section 3 describes each of the tools utilized in the I2M pipeline. Section 4 describes each of the above contributions in detail. Section 5.2 shows how the aforementioned I2M conversion requirements are satisfied by both the I2M pipeline and PODM. Section 5.3 shows the impact of the hierarchical load balancing model as it is tested on different supercomputers. Section 5.4 shows the performance improvement from the optimized local reconnection algorithm and the performance of both CDT3D and PODM when generating large-size meshes (up to about 100 million elements). Section 6 provides an analysis of our results while section 7 discusses needed future work and section 8 concludes our paper.
2 Related Work
@@
There are numerous methods that process medical data as input to create meshes. The type of meshes generated must be taken into account when considering the target application. Meshes can be classified into two categories - structured and unstructured. While some structured mesh generation methods have been shown to provide accurate results within CFD simulations [AcceleratedEstimation2021, DynamicOfBloodFlows2021, Enhanced4D2022], these types of methods can sometimes fail to capture features within complex geometries [SAMRBook1999]. Unstructured meshes can be further classified into two categories - isotropic and anisotropic. Isotropic mesh generation methods are well suited for high curvature geometries. Anisotropic meshes contain high aspect ratio elements that include directional information (e.g., hemodynamics). Additionally, they may contain fewer elements in some cases compared to isotropic meshes generated from the same image. A mesh should ideally be as small as possible to reduce the memory and CPU requirements for the solver. These benefits make anisotropic meshes suitable for blood flow problems.
@@
There are many methods that convert 3D images into 3D meshes. For a comprehensive overview of these methods, see [ZhangBook2016]. Many 3D isotropic methods utilize Delaunay refinement techniques [Boltcheva, CGAL2007, Pons2007, AnisotropicFluidSolidI2M2010, foteinos2014high, PDRPODMExperience]. A challenge with Delaunay refinement, which still remains an open problem, is that almost flat tetrahedra (slivers) can survive known heuristics designed to remove them [Chew:1997, cheng2000sliver, Boltcheva]. These low quality elements introduce error when processed by a solver. Additional I2M methods utilize lattice space-tree (octree) decomposition techniques [LatticeCleaving, ZhangAutomaticMultimaterial2010, Zhang3DFinite2005, LD_Chernikov, ZhangChallenges2013] to provide meshes with adaptively-sized, high quality elements. Some also utilize the Dual Contouring technique to improve element quality [DualContour3DSegmentationAnisotropic2018, ParallelDC2013]. Others warp the surfaces of the lattice to the boundaries of the object using either a mass-spring system, an FE constitutive model, or an optimization scheme [Molino03acrystalline, CBC3DYixun, Fuchs20011400]. Some methods focus on processing surface geometries of medical data to create structured Cartesian cut-cell meshes [SolutionAdaptive2021] and isotropic tetrahedral meshes [AssessmentMeshGen2008]. The study in [MeshGenNRR2008] evaluates several of these sequential isotropic-based methods’ suitability for surgical simulations and non-rigid registration applications. There are methods which utilize parallel mesh generation techniques (with either CPUs [ParallelDC2013, AcceleratedEstimation2021, DynamicOfBloodFlows2021, Enhanced4D2022] or GPUs [ParallelDC2013]). The methods in [AcceleratedEstimation2021, DynamicOfBloodFlows2021, Enhanced4D2022] all focus on generating Cartesian cut-cell meshes while [ParallelDC2013] focuses on generating isotropic tetrahedral meshes from biomolecular complexes (as opposed to 3D images). [DualContour3DSegmentationAnisotropic2018] utilizes an anisotropic Giaquinta–Hildebrandt operator-based geometric flow method to smooth mesh surfaces while preserving surface and volume features. However, this smoothing is applied to isotropic meshes.
@@
Some methods focus on converting medical images into anisotropic meshes. Although potentially useful for blood flow simulations, [ClericiAnisotropic2020] focuses on using an anisotropic mesh adaptation procedure to construct a high-fidelity virtual model of a fractured pelvis, and then to build a physical replica (by means of additive manufacturing techniques) to represent the anatomical object from a segmented image. The method in [AnisotropicFluidSolidI2M2010] generates high-quality adaptive hybrid meshes for fluid-solid geometries within cardiac environments. Following segmentation, surfaces between heart tissue and blood (and other non-cardiovascular tissue/fluids) are extracted and adapted. Anisotropic tetrahedra are generated for the interior of heart tissue, a prismatic boundary layer is generated for the interface between tissue and blood, and TetGen (an isotropic boundary-constrained Delaunay-based method [TetGen2015]) is used to tetrahedralize the remainder of the blood region. Boundary layers in meshes are useful when calculating metrics such as pressure or wall shear stress (WSS) caused by blood flow [CFDAneurysmRupture2019, AnisotropicFluidSolidI2M2010]. Although the method generated high-quality hybrid meshes, the Delaunay-based isotropic elements in fluid flow regions (such as in the coronary arteries) were identified as a drawback, with anisotropy suggested as an improvement. Additionally, Delaunay tetrahedralization limits the quality of the tetrahedral core for complex biological geometries [AnisotropicFluidSolidI2M2010] and can fail to terminate for certain cases [EwCCBC3D]. Another method was developed to model cerebral aneurysms for fluid-structure interaction simulations [AnisotropicCerebralAneurysms2013]. The surface mesh of a healthy blood vessel is deformed onto an aneurysm model, simulating the formation of the aneurysm. While the simulation results demonstrated the benefit of utilizing anisotropic meshes rather than isotropic meshes, the mesh generation method is sequential, as with all the aforementioned adaptive anisotropic methods.
Two approaches are proposed to help streamline the discretization of complex vascular geometries within the numerical modeling process. The first combines multiple software tools into a single pipeline to provide the following: (1) image-to-mesh conversion which satisfies the quality, fidelity, and smoothness ($C^{0}$) requirements, (2) the generation of a boundary layer grid over the high fidelity surface, (3) a parallel adaptive anisotropic meshing procedure which provides near real-time performance, and (4) robustness, which is satisfied by the pipeline’s ability to process segmented images and CAD models [EwCCDT3D]. To the best of our knowledge, there is no single method that provides all of the above. Although it is a Delaunay-based method, the second approach was previously tested to generate a mesh of reasonably good fidelity and quality when converting a segmented image of an aneurysm into an isotropic mesh [EwCCBC3D]. Consequently, we extend its functionality to not only adapt a mesh based on the curvature of the image boundary but to also approximate a user-defined sizing function to allow adaptivity within the volume when generating isotropic meshes (potentially making this method suitable for vascular flow simulations, where the next step is to test the method with error-based sizing criteria within a simulation). While some of the aforementioned methods are capable of generating adaptive (within the volume) isotropic meshes of good quality, these methods are sequential [CGAL2007, Pons2007, ZhangAutomaticMultimaterial2010, AnisotropicFluidSolidI2M2010, LatticeCleaving] (as opposed to our second method’s parallel speculative execution model that offers real-time performance even when generating large meshes). We test both techniques’ (converting segmented medical images to either isotropic or anisotropic meshes) feasibility as options to satisfy medical image-to-mesh conversion requirements when generating small-size meshes (e.g., 1 million elements) and much larger meshes (up to 100 million elements).
3 Background for Tools in Adaptive Anisotropic Image-to-mesh Conversion Pipeline
Details of each tool utilized in the I2M pipeline to generate adaptive anisotropic meshes are given. Section 3.1 describes the lattice-based I2M software CBC3D, which generates isotropic meshes of high quality, good fidelity, and smoothness. Section 3.2 describes the Advancing Front Local Reconnection (AFLR) software utilized to generate a boundary layer grid over the CBC3D-generated mesh surface. Section 3.3 gives details about the parallel adaptive anisotropic mesh generation software CDT3D.
3.1 Image-to-mesh Conversion